Consider a vector field F = (72 - y + z) - (2rry + y) i. Then f is given by: Select one: 2 0릉 콜 블 0 y 를 쁠 릉 t쁨 쁠
Answer #1

(4+2) - (224+y); Foc = (3-4+36) Fy = -(2xy+y) Fxc = да (F (3,4)) FC2;4) = (Ford $632-4*2) de - F(8,4) ser 23 aght sen 2y F (